OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! For what value of k, k > 0, does the integral from k to 0 of the quantity 3 times k minus 2 times k times x, dx equals k squared. a. 1 b. 2 c. 3 d. 4

OpenStudy (xguardians):

Can you please write out the equation, I can assist you from there.

OpenStudy (aryana_maria2323):

\[\int\limits_{k}^{0}(3k-2kx)dx=k^2\]

OpenStudy (aryana_maria2323):

@xGuardians

OpenStudy (andrewyates):

\[\int_k^0 (3k-2kx) \ dx = k^2 \\ = \int_k^0 3k \ dx - \int_k^0 2kx \ dx \\ = 3k\int_k^0 dx - 2k\int_k^0 x \ dx = k^2 \] Can you solve this equation for \(k\)?

OpenStudy (aryana_maria2323):

which part of the equation should I put into my calculator?

OpenStudy (andrewyates):

First, you need to evaluate the integrals.

OpenStudy (aryana_maria2323):

evaluate what integral?

OpenStudy (andrewyates):

Let's back up, do you understand how I got \[ 3k\int_k^0 dx - 2k\int_k^0 x \ dx = k^2 \]

OpenStudy (aryana_maria2323):

Yes I understand your steps above

OpenStudy (andrewyates):

Okay, so now we can go one step further and evaluate the integrals like so: \[ = 3k \bigg[ x \bigg]_k^0 - 2k \bigg[ \frac{x^2}{2} \bigg]_k^0 = k^2 \] Does this make sense?

OpenStudy (aryana_maria2323):

Yes it does

OpenStudy (aryana_maria2323):

@andrewyates

OpenStudy (andrewyates):

So now, \[ = 3k(0-k)-2k(\frac{0^2}{2} - \frac{k^2}{2}) \\ = -3k^2+k^3 \]

OpenStudy (andrewyates):

Which is equal to \(k^2\): \[ -3k^2+k^3 = k^2 \implies \\ k^3-4k^2 = 0 \implies \\ k^2(k-4) = 0 \implies \\ k = 0, 4 \]

OpenStudy (aryana_maria2323):

Okay but that is not one of my answers?

OpenStudy (aryana_maria2323):

@andrewyates

OpenStudy (andrewyates):

k = 4 is one of your answers. D.

OpenStudy (aryana_maria2323):

ohh haha I was going with 0,4